can you find probability of a if yiu have probability of b and probability of a intersect b
Answer #1

Answer)

Yes, probability of a, when b has already happened (P(A/B))

Is given by

P(A/B) = P(A intersect B)/(P(B)

It is written as,

P(A/B) = P(A and B)/P(B)

Where, P(A and B) is same as P(A intersect B).